Find the domain of the function.
(a) (2x+1)/(x^2+x-2)
(b) cube square of x/x^2+1
(c)(sqrt4-x)+(sqrtx^2 -1)

Answer by stanbon(75887)

Find the domain of the function.
(a) (2x+1)/(x^2+x-2)
Factor: (2x+1)/[(x+2)(x-1)
Domain: All Real Numbers except x = -2 and x = 1
Interval Notation: (-oo,-2]U(-2,1]U(1,+oo)
--------------------------------------
(b) cbrt[x/x^2+1]
Domain: All Real Numbers
Interval Notation: (-oo,+oo)
--------------------------------------
(c)(sqrt(4-x)+sqrt(x^2-1)
------
= sqrt(4-x) + sqrt[(x+1)(x-1)]
Solve (x+1)(x-1) >= 0
x <= -1 or x >= 1
Interval Notation: (-oo,-1]U[1,+oo)
-------------------
Solve 4-x >= 0
x <= 4
----
Domain: All Real Numbers x <=-1 or 1<= x <=4
Interval Notation: (-oo,-1]U[1,4]
==========================
Cheers,
Stan H.
==========================